Just Cabinet Manufacturing Inc owns machinery used to make cabinets. The original cost of the machinery was $20,000 and it was purchased July 1, 2014. Deprecation has been recorded yearly at $2,400. The current accumulated depreciation as of Dec 31, 2017 is $8,400. Just Cabinet sold the machinery on September 1, 2018 for $5,200 cash.

Prepare the journal entry to:

#1-update the deprecation for the year of the sale (2018)

#2-record the sale
Answer #1

Depreciation expense from January 1, 2018 to September 1, 2018 = 2,400 x 8/12

= $1,600

Accumulated depreciation till the date of sale = 8,400 + 1,600

= $10,000

Book value of machinery on the date of sale = 20,000 - 10,000

= $10,000

Sale price of machinery = $5,200

Loss on sale of machinery = Book value of machinery on the date of sale - Sale price of machinery

= 10,000 - 5,200

= $4,800

Journal

No

Account Title

Debit

Credit

1. Depreciation expense 1,600
Accumulated depreciation - Machinery 1,600
2. Cash 5,200
Accumulated depreciation - Machinery 10,000
Loss on sale of machinery 4,800
Machinery 20,000

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.